OpenStudy (anonymous):

each of my daughters has as many brothers as sisters, but each of my sons has three times as many sisters as brothers. how many children do i have?

OpenStudy (anonymous):

please help i am confused

OpenStudy (hoblos):

2 sons and 3daughters each daughter has 2 sisters and 2 brothers each son has 3 sisters and 1 brother

OpenStudy (hoblos):

to solve... let x= number of sons y=number of daughters so y-1=x and 3(x-1) =y solve the system

OpenStudy (anonymous):

Ripped off: Lets assume number of daughters = x and number of sons = y each daughter has as many sisters as brothers so, y = x - 1 x - y = 1 ----------- (1) each son has 3 times as many sisters as brothers so, x =3(y - 1) x = 3y - 3 x - 3y = -3 ---------------(2) Subtract (2) from (1), -y + 3y = 1 - (-3) 2y = 4 y = 4/2 = 2 From (1), x - y = 1 x - 2 = 1 x = 1 + 2 = 3 Daughters = x = 3 Sons = y = 2

OpenStudy (anonymous):

Okay I am too lazy now ( http://au.answers.yahoo.com/question/index?qid=20110901003913AAhDwM8)

OpenStudy (anonymous):

thank you all